OpenStudy (anonymous):

Given f(x) = the quantity of 2x minus 2 divided by 4, solve for f-1(3). 1 5 7 8 @AaronAndyson

OpenStudy (aaronandyson):

same as i had earlier explained.

OpenStudy (anonymous):

may you explain it again :( I get it when you tell me to plug this and that in and ill multiple and stuff. can you do that again please lol

OpenStudy (aaronandyson):

lol okay f(x) = the quantity of 2x minus 2 divided by 4, thats means f(x)= 2x-2/4 correct?

OpenStudy (anonymous):

yes

OpenStudy (aaronandyson):

f-1(3) so this is f(-3).

OpenStudy (aaronandyson):

so plug 3 in the x

OpenStudy (aaronandyson):

f(x) = 2(3) - 2/4 what's the answer?

OpenStudy (anonymous):

Aaron are you sire f-1 (3) does not represent the inverse function?

OpenStudy (aaronandyson):

i'm giving my best shot here...-.-

OpenStudy (anonymous):

2*3=6 6-2/4? im not so sure about that one

OpenStudy (anonymous):

OK but f-1 is the representation of an inverse function

OpenStudy (aaronandyson):

yes so whats 6-2?

OpenStudy (anonymous):

4

OpenStudy (aaronandyson):

so whats 4/4?

OpenStudy (anonymous):

oh so then I do 4/4=1

OpenStudy (aaronandyson):

yes.

OpenStudy (anonymous):

This is not correct. You need to find the inverse function for f(x) = (2x-2) / 4 first then solve.